# On semidefinite descriptions for convex hulls of quadratic programs

### 07 Mar 2024 v1

@@ 1 Introduction
where $q_{\textup{obj}},q_{1},\dots,q_{m}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ are each (possibly nonconvex) quadratic functions. For each $i\in\left\{{\textup{obj}}\right\}\cup[m]$, we will write $q_{i}(x)=x^{\top}A_{i}x+2b_{i}^{\top}x+c_{i}$ for $A_{i}\in{\mathbb{S}}^{n}$, $b_{i}\in{\mathbb{R}}^{n}$, and $c_{i}\in{\mathbb{R}}$. These optimization problems arise naturally in a variety of application areas (see e.g., [25] and references therein).
@@
Although QCQPs are NP-hard in general, they admit a natural tractable convex relaxation known as the standard (Shor) semidefinite program (SDP) relaxation,
@@
$\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{x\in{\mathbb{R}}^{n}}\left\{\left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}:\,\begin{array}[]{l}\exists X\succeq xx^{\top}:\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\end{array}\right\}.$
@@
A recent and growing line of work has offered deterministic conditions under which the SDP relaxation of a general QCQP is exact for various definitions of exactness. In their celebrated paper, Fradkov and Yakubovich [8] prove the S-lemma, which implies that the problem of minimizing an arbitrary quadratic objective function over the unit ball (or any single quadratic constraint) can be solved via SDP techniques. Specifically, the S-lemma implies that objective value exactness—the condition that the optimal value of the QCQP and the optimal value of its SDP relaxation coincide—holds for QCQPs with a single constraint. Locatelli [16], Jeyakumar and Li [14], Bomze et al. [4] present sufficient conditions for objective value exactness to continue to hold in the presence of additional linear constraints and/or one additional quadratic constraint (possibly using strengthened relaxations). Burer and Ye [6] study diagonal QCQPs—those QCQPs for which $A_{\textup{obj}},A_{1},\dots,A_{m}$ are diagonal matrices—with a general number of constraints and give sufficient conditions for objective value exactness. Sojoudi and Lavaei [21] study sparse SDPs where the $A_{i}$ matrices have an underying graph structure and give sufficient conditions for objective value exactness phrased in terms of the graph structure and the associated coefficients. Wang and Kılınç-Karzan [25] continue this line of work by developing a general framework for deriving sufficient conditions for both objective value exactness and convex hull exactness—the condition that the convex hull of the QCQP epigraph coincides with the (projected) SDP epigraph—for QCQPs where a specific dual set, $\Gamma_{1}$, is polyhedral (see Section 2). Beyond being a natural sufficient condition for objective value exactness, convex hull exactness has its own applications and motivation, particularly in the derivation of strong relaxations of certain critical substructures in nonconvex problems; see e.g., [23, 25] and references therein.
@@
While the framework presented in [25] at once covers and extends many existing results on exactness [15, 6, 8, 12, 18, 26, 5], it is still quite limited. In particular, the assumption that $\Gamma_{1}$ is polyhedral is rarely satisfied outside of simultaneously diagonalizable QCQPs and precludes the results in [25] from being applicable to a wider range of interesting QCQPs. In this paper, we generalize the results introduced in [25] by eliminating its reliance on the polyhedrality assumption. Specifically, after formally defining our setup and recalling basics on the SDP relaxation of a QCQP, and the cone of convex Lagrange multipliers $\Gamma$ (and the closely related sets $\Gamma_{1}$ and $\Gamma^{\circ}$) in Section 2, we present in Section 3 a sufficient condition for convex hull exactness that generalizes [25, Theorem 1]. This sufficient condition (Theorem 1) states that convex hull exactness holds as long as certain systems of equations (that depend on $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$) contain nontrivial solutions. Thus, this sufficient condition can be checked in a systematic manner whenever $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$ is sufficiently simple. Furthermore, in Theorem 2, we show that our sufficient condition for convex hull exactness is also necessary under the assumption that $\Gamma^{\circ}$ is facially exposed (see Assumption 2 and its surrounding discussion). To the best of our knowledge, this is the first necessary and sufficient condition for convex hull exactness even in the context of diagonal QCQPs (where $\Gamma$ is automatically polyhedral). Based on these results, we also revisit the previous polyhedral setting of [25] and provide more refined necessary and sufficient conditions for convex hull exactness in this setting (see Sections 3.4 and 5). These results offer an algebraic way to verify convex hull exactness even when the QCQP is not simultaneously diagonalizable. We illustrate our results (Theorems 1 and 3.4) on three applications (a prototypical mixed binary set arising in sparse regression problems, quadratic matrix programs (QMPs), and the sets defined by two quadratic constraints) in Appendix B. In the first two applications, the resulting $\Gamma$ sets are non-polyhedral, and thus the sufficient conditions from [25] that work under the polyhedrality assumption of $\Gamma$ are not applicable. For QMPs, we show that convex hull exactness holds whenever the number of constraints is small (when compared to the rank of the matrix variable) which strengthens the objective value exactness result of [2]. In the last application, we illustrate that our Section 3.4 reduces to a succinct and easy to check necessary and sufficient condition for convex hull exactness when $q_{\textup{obj}}=0$ and $m=2$ as opposed to the more difficult to check condition of [26].
@@ 1.1 Notation
For a positive integer $n$, let $[n]\coloneqq\left\{1,\dots,n\right\}$. For $i\in[n]$, let $e_{i}$ be the $i$th standard basis vector in ${\mathbb{R}}^{n}$. Let $0_{n}$ denote the zero vector in ${\mathbb{R}}^{n}$. Let ${\mathbb{S}}^{n}$ denote the vector space of $n\times n$ real symmetric matrices and ${\mathbb{S}}^{n}_{+}$ the cone of positive semidefinite matrices. For $M\in{\mathbb{S}}^{n}$, we write $M\succeq 0$ (resp. $M\succ 0$) to denote that $M$ is positive semidefinite (resp. positive definite). Let $\ker(M)$ denote the kernel of $M$. Suppose ${\cal M}$ is a subset of some Euclidean space. Let $\operatorname{conv}({\cal M})$, $\operatorname{clcone}({\cal M})$, $\operatorname{span}({\cal M})$, ${\cal M}^{\perp}$, $\operatorname{rint}({\cal M})$, and $\dim({\cal M})$ denote the convex hull, closed conic hull, span (linear hull), orthogonal complement, relative interior, and dimension of ${\cal M}$ respectively. Specifically, if $\operatorname*{\mathbb{E}}$ is the Euclidean space containing ${\cal M}$, then ${\cal M}^{\perp}\coloneqq\left\{x\in\operatorname*{\mathbb{E}}:\,\left\langle x,m\right\rangle=0,\,\forall m\in{\cal M}\right\}$. Suppose $K$ is a cone in some Euclidean space. Let $K^{\circ}$ denote the polar cone of $K$. The notation $F\trianglelefteq K$ denotes that $F$ is a face of $K$. By convention, faces of cones are always nonempty. Given $x,y$ in some Euclidean space $\operatorname*{\mathbb{E}}$, let $[x\pm y]\coloneqq\operatorname{conv}(\left\{x-y,x+y\right\})$.
@@ 2 Preliminaries
We study the epigraph of the QCQP in (1) defined as
@@
It is well known in the QCQP literature [3, 25, 10] that the SDP relaxation of a QCQP is equivalent (under standard assumptions) to the dual of the Lagrangian dual. We start by introducing notation related to Lagrangian aggregation and then state this equivalence formally in Remark 1.
Let $q:{\mathbb{R}}^{n}\to{\mathbb{R}}^{1+m}$ be indexed by $\left\{{\textup{obj}}\right\}\cup[m]$ where $q(x)_{\textup{obj}}=q_{\textup{obj}}(x)$ and $q(x)_{i}=q_{i}(x)$ for $i\in[m]$. Let $e_{\textup{obj}},e_{1},\dots,e_{m}$ denote the corresponding unit vectors in ${\mathbb{R}}^{1+m}$. Note that for $(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{1+m}$, the expression $\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle=\gamma_{\textup{obj}}q_{\textup{obj}}(x)+\sum_{i=1}^{m}\gamma_{i}q_{i}(x)$ is a quadratic function in $x\in{\mathbb{R}}^{n}$. Define $A(\gamma_{\textup{obj}},\gamma)\coloneqq\gamma_{\textup{obj}}A_{\textup{obj}}+\sum_{i\in[m]}\gamma_{i}A_{i}$. Similarly, define $b(\gamma_{\textup{obj}},\gamma)$, and $c(\gamma_{\textup{obj}},\gamma)$.
@@
We recall and extend the following definition from [25].
@@
We define $\Gamma$ as the cone of convex Lagrange multipliers for (1), i.e.,
$\Gamma\coloneqq\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}_{+}\times{\mathbb{R}}^{m}_{+}:\,A(\gamma_{\textup{obj}},\gamma)\succeq 0\right\},\qquad\text{and also its cross section}$
$\Gamma_{1}\coloneqq\left\{\gamma\in{\mathbb{R}}_{+}^{m}:\,(1,\gamma)\in\Gamma\right\}=\left\{\gamma\in{\mathbb{R}}_{+}^{m}:\,A[\gamma]\succeq 0\right\}.$
These sets are related to the feasible domain of a partial dual of the SDP relaxation (see Remark 2). Note that given $(\gamma_{\textup{obj}},\gamma)\in\Gamma$, the quadratic function $x\mapsto\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle$ is convex. Similarly, given $\gamma\in\Gamma_{1}$, the quadratic function $x\mapsto[\gamma,q(x)]$ is convex.
We will make the following blanket assumption for the remainder of the paper. This assumption can be interpreted as a dual strict feasibility condition and is standard in the literature [6, 25].
@@
Assumption 1 is equivalent to the assumption that there exists $\gamma^{*}\in\Gamma_{1}$ such that $A[\gamma^{*}]\succ 0$. One direction is immediate. In the other direction, suppose $A(\gamma^{*}_{\textup{obj}},\gamma^{*})\succ 0$. As the minimum eigenvalue of $A(\gamma^{*}_{\textup{obj}},\gamma^{*})$ is continuous in its arguments, we may assume that $\gamma^{*}_{\textup{obj}}>0$. Then, $A[\gamma^{*}/\gamma^{*}_{\textup{obj}}]=\frac{1}{\gamma^{*}_{\textup{obj}}}A(\gamma^{*}_{\textup{obj}},\gamma^{*})\succ 0$.
@@
Under Assumption 1, $\Gamma$ is the closed cone generated by its slice at $\gamma_{\textup{obj}}=1$, i.e., $\Gamma=\operatorname{clcone}\left(\left\{(1,\gamma):\,\gamma\in\Gamma_{1}\right\}\right)$. (See discussion following [25, Assumption 2].)
@@
${\cal S}_{\textup{SDP}}\coloneqq\left\{(x,t)\in{\mathbb{R}}^{n}\times{\mathbb{R}}:\,\begin{array}[]{l}\exists X\succeq xx^{\top}:\\ \left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}\leq 2t\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\end{array}\right\},$ (5)
and $\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{(x,t)\in{\cal S}_{\textup{SDP}}}2t$. By taking $X=xx^{\top}$ in (5), we see that $\operatorname{Opt}\geq\operatorname{Opt}_{\textup{SDP}}$ and $\operatorname{conv}({\cal S})\subseteq{\cal S}_{\textup{SDP}}$.
The following lemma rewrites ${\cal S}_{\textup{SDP}}$ in terms of $\Gamma$, $\Gamma_{1}$, and $\Gamma^{\circ}$ and follows from a straightforward duality argument; see Figure 1 for an illustration of these sets. These expressions allow for a straightforward analysis of ${\cal S}_{\textup{SDP}}$ whenever the corresponding dual set is sufficiently simple. {restatable}lemmalemSSDP Suppose Assumption 1 holds. Then,
@@
=$=\left\{(x,t)\in{\mathbb{R}}^{n+1}:\begin{array}[]{r}\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle\leq 2\gamma_{\textup{obj}}t,\,\forall(\gamma_{\textup{obj}},\gamma)\in\Gamma\end{array}\right\}$
@@
The proof of Remark 1 is given in Appendix A. We list some corollaries of Remark 1.
@@
One may compute $\operatorname{Opt}_{\textup{SDP}}$ in the form (5) by minimizing over $x\in{\mathbb{R}}^{n}$ the value of an inner minimization problem over the matrix variable $X$. Writing $X=xx^{\top}+\xi$ and taking the SDP dual in the $\xi$ variable then results in the same saddle-point structure observed in Corollary 1. In other words, $\Gamma_{1}$ is the feasible domain to this partial dual of (5).
[…]
Figure 1: The sets ${\cal S}$, ${\cal S}_{\textup{SDP}}$, $\Gamma$, and $\Gamma^{\circ}$ are shown in blue, green, orange, and yellow respectively. By Remark 1, $(x,t)\in{\cal S}_{\textup{SDP}}$ if and only if $q(x)-2te_{\textup{obj}}\in\Gamma^{\circ}$.
Our results and analysis depend on the geometry of $\Gamma$, $\Gamma^{\circ}$, and their faces. We close this section with some properties of convex cones and their faces specialized to our setting. See [1] for more general statements.
@@
Similarly, define the conjugate face of ${\cal F}$ for a face ${\cal F}\trianglelefteq\Gamma$.
@@
Given $(x,t)\in{\cal S}_{\textup{SDP}}$, we have that $q(x)-2te_{\textup{obj}}\in\operatorname{rint}({\cal G}(x,t))$ and ${\cal F}(x,t)=\Gamma\cap(q(x)-2te_{\textup{obj}})^{\perp}$.

### 20 Mar 2024 v2

@@ 1 Introduction
where $q_{\textup{obj}},q_{1},\dots,q_{m}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ are each (possibly nonconvex) quadratic functions. For each $i\in[0,m]$, we will write $q_{i}(x)=x^{\top}A_{i}x+2b_{i}^{\top}x+c_{i}$ for $A_{i}\in{\mathbb{S}}^{n}$, $b_{i}\in{\mathbb{R}}^{n}$, and $c_{i}\in{\mathbb{R}}$. These optimization problems arise naturally in applications (see e.g., [18] and references therein).
@@
Although QCQPs are NP-hard in general, they admit a natural convex relaxation known as the standard semidefinite program (SDP) relaxation,
@@
$\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{\begin{subarray}{c}x\in{\mathbb{R}}^{n}\\ X\in{\mathbb{S}}^{n}\end{subarray}}\left\{\left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}:\,\begin{array}[]{l}\left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\\ X\succeq xx^{\intercal}\end{array}\right\}.$
@@
A growing line of work has offered deterministic conditions under which the SDP relaxation of a general QCQP is exact for various definitions of exactness. In their celebrated paper, Fradkov and Yakubovich [7] prove the S-lemma, which implies objective value exactness—the condition that the optimal value of the QCQP and the optimal value of its SDP relaxation coincide—holds for QCQPs with a single constraint. Locatelli [11], Jeyakumar and Li [10], Bomze et al. [4] present sufficient conditions for the same condition to hold under additional linear constraints and/or one additional quadratic constraint (possibly using strengthened relaxations). Burer and Ye [6] give sufficient conditions for objective value exactness for general diagonal QCQPs—those QCQPs for which $A_{\textup{obj}},A_{1},\dots,A_{m}$ are diagonal matrices. Sojoudi and Lavaei [16] study sparse SDPs where the $A_{i}$ matrices have an underying graph structure and give sufficient conditions for objective value exactness phrased in terms of the graph structure and the associated coefficients. Wang and Kılınç-Karzan [18] continue this line of work by developing a general framework for deriving sufficient conditions for both objective value exactness and convex hull exactness—the condition that the convex hull of the QCQP epigraph coincides with the (projected) SDP epigraph—for QCQPs where a specific dual set, $\Gamma_{1}$, is polyhedral (see section 2). Beyond being a natural sufficient condition for objective value exactness, convex hull exactness has its own applications and motivation, particularly in the derivation of strong relaxations of certain critical substructures in nonconvex problems; see e.g., [18] and references therein.
@@
While the framework presented in [18] at once covers many existing results on exactness [6, 7, 9, 13, 19, 5], it is still quite limited. In particular, the assumption that $\Gamma_{1}$ is polyhedral is rarely satisfied outside of simultaneously diagonalizable QCQPs and precludes the results in [18] from being applicable to a wider range of interesting QCQPs. In this paper, we generalize the framework of [18] by eliminating its reliance on the polyhedrality assumption. Specifically, theorem 1 states that convex hull exactness holds as long as certain systems of equations (that depend on $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$) contain nontrivial solutions. This sufficient condition can be checked systematically if $\Gamma$, $\Gamma_{1}$, or $\Gamma^{\circ}$ is sufficiently simple. Furthermore, theorem 2 states that this sufficient condition is also necessary under the assumption that $\Gamma^{\circ}$ is facially exposed (see Assumption 2). To the best of our knowledge, this is the first necessary and sufficient condition for convex hull exactness even in the context of diagonal QCQPs (where $\Gamma$ is automatically polyhedral and facially exposed). In section 3.4, we provide refined necessary and sufficient conditions for convex hull exactness in the previous polyhedral setting of [18]. We illustrate our results (theorems 1 and 3.4) on two applications (quadratic matrix programs (QMPs) and sets defined by two quadratic constraints) in section 4. In the first application, the resulting $\Gamma$ set is non-polyhedral, and thus the sufficient conditions from [18] are not applicable. In the second application, we illustrate that our section 3.4 reduces to a succinct and easy to check necessary and sufficient condition for convex hull exactness when $q_{\textup{obj}}=0$ and $m=2$.
@@ 1.1 Notation
For a positive integer $n$, let $[n]\coloneqq\left\{1,\dots,n\right\}$ and let $[0,n]\coloneqq\left\{0,1,\dots,n\right\}$. For $i\in[n]$, let $e_{i}$ be the $i$th standard basis vector in ${\mathbb{R}}^{n}$. Let $0_{n}$ denote the zero vector in ${\mathbb{R}}^{n}$. Let ${\mathbb{S}}^{n}$ denote the set of $n\times n$ real symmetric matrices and ${\mathbb{S}}^{n}_{+}$ the cone of positive semidefinite (PSD) matrices. For $M\in{\mathbb{S}}^{n}$, we write $M\succeq 0$ (resp. $M\succ 0$) to denote that $M$ is PSD (resp. positive definite). Let $\ker(M)$ denote the kernel of $M$. Suppose ${\cal M}$ is a subset of some Euclidean space. Let $\operatorname{conv}({\cal M})$, $\operatorname{clcone}({\cal M})$, $\operatorname{span}({\cal M})$, ${\cal M}^{\perp}$, $\operatorname{rint}({\cal M})$, and $\dim({\cal M})$ denote the convex hull, closed conic hull, span (linear hull), orthogonal complement, relative interior, and dimension of ${\cal M}$ respectively. Specifically, if $\operatorname*{\mathbb{E}}$ is the Euclidean space containing ${\cal M}$, then ${\cal M}^{\perp}\coloneqq\left\{x\in\operatorname*{\mathbb{E}}:\,\left\langle x,m\right\rangle=0,\,\forall m\in{\cal M}\right\}$. Suppose $K$ is a cone in some Euclidean space. Let $K^{\circ}$ denote the polar cone of $K$. The notation $F\trianglelefteq K$ denotes that $F$ is a face of $K$. By convention, faces of cones are always nonempty. Given $x,y$ in some Euclidean space $\operatorname*{\mathbb{E}}$, let $[x\pm y]\coloneqq\operatorname{conv}(\left\{x-y,x+y\right\})$.
@@ 2 Preliminaries
We study the epigraph of the QCQP in (1):
@@
It is well known in the QCQP literature [3, 18, 8] that the SDP relaxation of a QCQP is equivalent (under standard assumptions) to the dual of the Lagrangian dual. We introduce notation related to Lagrangian aggregation and then state this equivalence formally in remark 1.
Let $q:{\mathbb{R}}^{n}\to{\mathbb{R}}^{1+m}$ be indexed by $[0,m]$ where $q(x)_{i}=q_{i}(x)$ for $i\in[0,m]$. Let $e_{\textup{obj}},e_{1},\dots,e_{m}$ denote the corresponding unit vectors in ${\mathbb{R}}^{1+m}$. Note that for $(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{1+m}$, the expression $\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle=\sum_{i=0}^{m}\gamma_{i}q_{i}(x)$ is a quadratic function in $x\in{\mathbb{R}}^{n}$. Define $A(\gamma_{\textup{obj}},\gamma)\coloneqq\sum_{i=0}^{m}\gamma_{i}A_{i}$. Similarly, define $b(\gamma_{\textup{obj}},\gamma)$, and $c(\gamma_{\textup{obj}},\gamma)$.
@@
We extend the following definition from [18].
@@
The cone of convex Lagrange multipliers for (1) and its cross section at $\gamma_{0}=1$ are respectively
$\Gamma\coloneqq\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}_{+}\times{\mathbb{R}}^{m}_{+}:\,A(\gamma_{\textup{obj}},\gamma)\succeq 0\right\}$
$\Gamma_{1}\coloneqq\left\{\gamma\in{\mathbb{R}}_{+}^{m}:\,A[\gamma]\succeq 0\right\}.$
These sets are related to the feasible domain of a partial dual of the SDP relaxation (see remark 2). Note that $\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle$ is a convex function in $x$ for any $(\gamma_{\textup{obj}},\gamma)\in\Gamma$. Similarly, $[\gamma,q(x)]$ is convex in $x$ for any $\gamma\in\Gamma_{1}$.
We make the following assumption for the remainder of the paper. This assumption can be interpreted as a dual strict feasibility condition and is standard [6, 18].
@@
Assumption 1 is equivalent to the assumption that there exists $\gamma^{*}\in\Gamma_{1}$ such that $A[\gamma^{*}]\succ 0$.
@@
Under Assumption 1, it holds that $\Gamma=\operatorname{clcone}\left(\left\{(1,\gamma):\,\gamma\in\Gamma_{1}\right\}\right)$. (See discussion following [18, Assumption 2].)
@@
$\begin{gathered}{\cal S}_{\textup{SDP}}\coloneqq\left\{(x,t)\in{\mathbb{R}}^{n}\times{\mathbb{R}}:\,\begin{array}[]{l}\exists X\succeq xx^{\top}:\\ \left\langle A_{\textup{obj}},X\right\rangle+2b_{\textup{obj}}^{\top}x+c_{\textup{obj}}\leq 2t\\ \left\langle A_{i},X\right\rangle+2b_{i}^{\top}x+c_{i}\leq 0,\,\forall i\in[m]\end{array}\right\},\qquad\operatorname{Opt}_{\textup{SDP}}\coloneqq\inf_{(x,t)\in{\cal S}_{\textup{SDP}}}2t\end{gathered}$ (6)
Taking $X=xx^{\top}$ in (6) gives $\operatorname{Opt}\geq\operatorname{Opt}_{\textup{SDP}}$ and $\operatorname{conv}({\cal S})\subseteq{\cal S}_{\textup{SDP}}$.
The following lemma rewrites ${\cal S}_{\textup{SDP}}$ in terms of $\Gamma$, $\Gamma_{1}$, and $\Gamma^{\circ}$ and follows from a straightforward duality argument. These expressions allow for a straightforward analysis of ${\cal S}_{\textup{SDP}}$ if the corresponding dual set is sufficiently simple. {restatable}lemmalemSSDP Suppose Assumption 1 holds. Then,
@@
=$=\left\{(x,t)\in{\mathbb{R}}^{n+1}:\,\left\langle(\gamma_{\textup{obj}},\gamma),q(x)\right\rangle\leq 2\gamma_{\textup{obj}}t,\,\forall(\gamma_{\textup{obj}},\gamma)\in\Gamma\right\}$
@@
Proof.
Fix $(x,t)\in{\mathbb{R}}^{n+1}$. Note that
$\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]=\inf_{\xi\in{\mathbb{S}}^{n}_{+}}\left\{q_{\textup{obj}}(x)+\left\langle A_{\textup{obj}},\xi\right\rangle:q_{i}(x)+\left\langle A_{i},\xi\right\rangle\leq 0,\,\forall i\in[m]\right\}.$
@@
We can interpret $\operatorname{Opt}_{\textup{SDP}}$ in (6) as minimizing over $x\in{\mathbb{R}}^{n}$ the value of an inner minimization problem over the matrix variable $X$. Writing $X=xx^{\top}+\xi$ and taking the SDP dual in $\xi$ results in the same saddle-point structure observed in corollary 1. In other words, $\Gamma_{1}$ is the feasible domain to this partial dual of (6).
Our results and analysis depend on the geometry of $\Gamma$, $\Gamma^{\circ}$, and their faces. We recall properties of convex cones and their faces specialized to our setting. See [1] for a general reference.
@@
Similarly, define the conjugate face for ${\cal F}\trianglelefteq\Gamma$.
@@
For $(x,t)\in{\cal S}_{\textup{SDP}}$, we have $q(x)-2te_{\textup{obj}}\in\operatorname{rint}({\cal G}(x,t))$ and ${\cal F}(x,t)=\Gamma\cap(q(x)-2te_{\textup{obj}})^{\perp}$.